OpenStudy (anonymous):

HELP PLEASE! The first and last terms of a 10-term arithmetic series are listed in the table. What is the sum of the series? Term Number Term 1 3 10 75 195 390 780 1125

undeadknight26 (undeadknight26):

Hmm can you supply me with maybe a screen shot of the graph? might help me a bit...

OpenStudy (***[isuru]***):

\[S _{n} = \frac{ n }{ 2} \times ( a + l )\]

OpenStudy (***[isuru]***):

l is the last term

OpenStudy (***[isuru]***):

\[S _{10} = \frac{ 10 }{ 2 } ( 3 + 75)\] wt will u get then ?

OpenStudy (anonymous):

390 :D thank you again @***[ISURU]***